# On quantum learning algorithms for noisy linear problems

### v1

@@ 2.3.3 Performance
3 An RLWE algorithm with quantum samples
@@
In this section it will be shown that the RLWE problem with quantum samples is reduced to the LWE problem with quantum samples by efficiently transforming a RLWE sample to an LWE sample. The reduction utilizes linearity of operations in $\mathcal{R}$. To be more specific, a RLWE sample is transformed to an LWE sample in a way that coefficients of a polynomial becomes a vector. The LWE algorithm described in Section 2.3 is then applied to recover $s(X)$. Grilo et al. mentioned in their paper [12] that the LWE algorithm is not applicable to the RLWE problem due to the absence of a proper isomorphism from $\mathcal{R}_{q}$ to $\mathbb{Z}_{q}^{n}$. Here we show that it can be done by employing the linearity.
3.0.1 Quantum sample
@@
Let ${\bm{s}}(X)$ be a secret polynomial and let $\mathcal{V}\subseteq\mathcal{R}_{q}$. A quantum RLWE sample for ${\bm{s}}$ and $\mathcal{V}$, denoted by $\lvert\psi^{\textup{rlwe}}_{{\bm{s}},\mathcal{V}}\rangle$ is defined as follows:
@@
$\lvert\psi_{s,\mathcal{V}}^{\textup{rlwe}}\rangle$ $=$ $\frac{1}{\sqrt{|\mathcal{V}|}}\sum_{a(X)\in\mathcal{V}}\lvert a(X)\rangle\lvert s(X)\!\cdot\!a(X)+e_{a}(X)\bmod q\rangle,\quad e_{a}(X)\leftarrow\chi_{\textup{rlwe}}\enspace$
@@
where $\chi_{\textup{rlwe}}$ is a probability distribution on $\mathcal{R}$ which samples $e_{a}(X)=\sum_{i=0}^{n-1}e_{a,i}X^{i}$ as $e_{a,0},\ldots,e_{a,n-1}\stackrel{{\scriptstyle i.i.d.}}{{\longleftarrow}}\chi$ using a probability distribution $\chi$ on $\mathbb{Z}$.
3.0.2 Transform from $\mathcal{R}_{q}$ to $\mathbb{Z}_{q}^{n}$
@@
Recall that the polynomial ring $\mathcal{R}$ is defined as $\mathcal{R}=\mathbb{Z}[X]/\langle\phi(X)\rangle$. Let us consider a (row) vector and a matrix constructed from a polynomial $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}$ as follows:
@@
$\mathsf{V}(a)=(a_{0},a_{1},\ldots,a_{n-1})\in\mathbb{Z}^{n},\quad\mathsf{M}_{\phi}(a)=\left(\begin{array}[]{ccc}&\mathsf{V}(a)&\\ &\mathsf{V}(X\!\cdot\!a\bmod\phi)&\\ &\vdots&\\ &\mathsf{V}(X^{n-1}\!\cdot\!a\bmod\phi)&\end{array}\right)\in\mathbb{Z}^{n\times n}$
@@
$\mathsf{V}(a\cdot b\bmod\phi)$ $=$ $\mathsf{V}(a)\,\mathsf{M}_{\phi}(b)\,,$
$\mathsf{V}(X\cdot a\bmod\phi)$ $=$ $\mathsf{V}(a)\,\mathbf{P}\,.$
Proof
@@
$a(X)\cdot b(X)\bmod\phi=\sum_{i=0}^{n-1}a_{i}X^{i}b(X)\bmod\phi=\sum_{i=0}^{n-1}a_{i}\big{(}X^{i}b(X)\bmod\phi\big{)}$
@@
Next, from $\phi(X)=X^{n}-\sum_{i=0}^{n-1}\phi_{i}X^{i}$ we have $X^{n}\equiv\sum_{i=0}^{n-1}\phi_{i}X^{i}\bmod\phi$, and thus
@@
$\mathsf{V}(X\cdot a\bmod\phi)$
$=$ $(0,a_{0},\ldots,a_{n-2})+a_{n-1}(\phi_{0},\phi_{1},\ldots,\phi_{n-1})$
@@
$=$ $\mathsf{V}(a)\,\mathbf{P}\enspace.$
Before continuing, let us introduce some notations. Let $\mathbf{M}^{T}$ denote the transpose of $\mathbf{M}\in\mathbb{R}^{l\times m}$. Let $(j\!+\!1)$-th column vector of $\mathbf{M}\in\mathbb{R}^{l\times m}$ be denoted by $\mathbf{M}_{j}\in\mathbb{R}^{l\times 1}$ so that $\mathbf{M}$ can be written as $\mathbf{M}=\left(\mathbf{M}_{0}\,|\,\cdots\,|\,\mathbf{M}_{m-1}\right)$. In particular, when $l=1$, $\mathbf{M}_{j}$ denotes the $(j\!+\!1)$-th coordinate of the row vector $\mathbf{M}$.
From Lemma 1, the matrix $\mathsf{M}_{\phi}(b)$ can be written as
@@
$(\mathsf{M}_{\phi}(b)_{j})^{T}$ $=$ $\big{(}\mathsf{V}(b)_{j},(\mathsf{V}(b)\mathbf{P})_{j},\ldots,(\mathsf{V}(b)\mathbf{P}^{n-1})_{j}\big{)}$ (8)
$=$ ($\mathsf{V}(b),((\mathbf{I}_{n})_{j})^{T}\right>,\left<\mathsf{V}(b),(\mathbf{P}_{j})^{T}\right>\ldots,\left<\mathsf{V}(b),((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}\,.$
3.0.3 RLWE to LWE
Let ${\bm{a}}=\mathsf{V}(a)$ for $a\in\mathcal{R}$. Define a unitary operation,
$U_{\phi,j}:\lvert{\bm{a}}\rangle\lvert 0\rangle\mapsto\lvert{\bm{a}}\rangle\lvert{\bm{a}}_{j}\rangle,\quad j\in\{0,\ldots,n-1\}$

### v2

@@ Introduction
Algorithm 2 $\mathsf{qLWE}\textup{-}\mathsf{Solver}$
1:oracle access to $\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$
2:$\tilde{{\bm{s}}}\in\mathbb{Z}_{q}^{n}$ or $\bot$
3:$i\leftarrow 0$
4:repeat
5:     $\big{|}\psi^{\textup{\tiny lwe}}_{{\bm{s}},V}\big{\rangle}\leftarrow\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$triangleright oracle query
6:     $\tilde{{\bm{s}}}\leftarrow\mathsf{BV}_{q,n}\big{(}\big{|}\psi^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}\big{\rangle}\big{)}$
7:     if  $\tilde{{\bm{s}}}=\bot$  then
8:         return $\bot$
9:     end if
10:     if  $\mathsf{Test}(\tilde{{\bm{s}}},\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}})=1$ then
11:         return $\tilde{{\bm{s}}}$
12:     end if
13:     $i\leftarrow i+1$
14:until $i\geq\ell_{1}$
15:return $\bot$
Algorithm 3 $\mathsf{Test}$
1:$\tilde{{\bm{s}}}\in\mathbb{Z}_{q}^{n}$, oracle access to $\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$
2:$1$ (pass) or 00 (fail)
3:$i\leftarrow 0$
4:repeat
5:     $({\bm{a}},b)\leftarrow\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$triangleright $b=\left<{\bm{a}},{\bm{s}}\right>+e_{\bm{a}}\bmod q$
6:     if $|b-\left<{\bm{a}},\tilde{{\bm{s}}}\right>\bmod q|>k$ then
7:         return 00
8:     end if
9:     $i\leftarrow i+1$
10:until $i\geq\ell_{2}$
11:return $1$
Revisiting previous works
This section presents new results. The first finding is about the solvability of the RLWE problem by reducing it to LWE. The second and third findings involve the development of efficient classical algorithms for SIS and size-reduced LWE problems under certain assumptions.
The purpose of the second and the third results is to show the following: if the assumptions on quantum samples taken in the previous works do hold, and then there also exist polynomial-time classical algorithms. A handful of classical algorithms are introduced below, but before presenting the results, we are urged to clarify the meaning of the word ‘classical’ here. In Result 2 and Result 3, we begin with quantum samples assumed in respective previous works [6, 8]. We do not question or evaluate the practicality of preparing quantum samples, but rather immediately carry out the measurement in computational basis. Each measurement gives rise to one classical sample that is put as an input to classical algorithms we develop. Therefore in estimating the complexity of classical algorithms, the number of samples used should be read as the number of quantum samples assumed.
Result 1: RLWE algorithm with quantum samples
@@
The goal of this subsection is to solve the RLWE problem given that the LWE problem is solvable by Algorithm 2. Applying Algorithm 2 to RLWE problem was originally unsuccessful by the authors of [6, Section V]. Below the RLWE problem with quantum samples is reduced to LWE by efficiently transforming a RLWE sample to an LWE sample. The reduction utilizes linearity of operations in $\mathcal{R}$. To be more specific, a RLWE sample is transformed to an LWE sample in a way that the coefficients of a polynomial become a vector. The LWE algorithm described in the previous section is then applied to recover $s(X)$.
@@
Let ${\bm{s}}(X)$ be a secret polynomial and let $\mathcal{V}\subseteq\mathcal{R}_{q}$. A quantum RLWE sample for ${\bm{s}}$ and $\mathcal{V}$, denoted by $\left|\psi^{\textup{\tiny rlwe}}_{{\bm{s}},\mathcal{V}}\,\right\rangle$ is defined as follows:
@@
$\left|\psi_{s,\mathcal{V}}^{\textup{\tiny rlwe}}\,\right\rangle=\frac{1}{\sqrt{|\mathcal{V}|}}\sum_{a(X)\in\mathcal{V}}\left|a(X)\,\right\rangle\left|s(X)\!\cdot\!a(X)+e_{a}(X)\bmod q\,\right\rangle,$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ e_{a}(X)\leftarrow\chi_{\textup{\tiny rlwe}}\enspace,$
@@
where $\chi_{\textup{\tiny rlwe}}$ is a probability distribution on $\mathcal{R}$ which samples $e_{a}(X)=\sum_{i=0}^{n-1}e_{a,i}X^{i}$ as $e_{a,0},\ldots,e_{a,n-1}\stackrel{{\scriptstyle i.i.d.}}{{\longleftarrow}}\chi$ using a probability distribution $\chi$ on $\mathbb{Z}$.
@@
Transform from $\mathcal{R}_{q}$ to $\mathbb{Z}_{q}^{n}$. Recall that the polynomial ring $\mathcal{R}$ is defined as $\mathcal{R}=\mathbb{Z}[X]/\langle\phi(X)\rangle$. Let us consider a (row) vector and a matrix constructed from a polynomial $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}$ as follows:
$\mathsf{V}(a)$ $=(a_{0},a_{1},\ldots,a_{n-1})\in\mathbb{Z}^{n},$
@@
$\mathsf{M}_{\phi}(a)$ $=\left(\begin{array}[]{ccc}&\mathsf{V}(a)&\\ &\mathsf{V}(X\cdot a\bmod\phi)&\\ &\vdots&\\ &\mathsf{V}(X^{n-1}\cdot a\bmod\phi)&\end{array}\right)\in\mathbb{Z}^{n\times n}\enspace,$
@@
$\mathsf{V}(a\cdot b\bmod\phi)=\mathsf{V}(a)\,\mathsf{M}_{\phi}(b)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathsf{V}(X\cdot a\bmod\phi)=\mathsf{V}(a)\,\mathbf{P}\enspace.$
Proof.
@@
$a(X)\cdot b(X)\bmod\phi=\sum_{i=0}^{n-1}a_{i}X^{i}b(X)\bmod\phi=\sum_{i=0}^{n-1}a_{i}\big{(}X^{i}b(X)\bmod\phi\big{)}\enspace,$
@@
Next, from $\phi(X)=X^{n}-\sum_{i=0}^{n-1}\phi_{i}X^{i}$, we have $X^{n}\equiv\sum_{i=0}^{n-1}\phi_{i}X^{i}\bmod\phi$, and thus
@@
$\phantom{=}\mathsf{V}(X\cdot a\bmod\phi)=(0,a_{0},\ldots,a_{n-2})+a_{n-1}(\phi_{0},\phi_{1},\ldots,\phi_{n-1})$
@@
$=\mathsf{V}(a)\,\mathbf{P}\enspace,$
the second equation holds. ∎
Let $\mathbf{M}^{T}$ denote the transpose of $\mathbf{M}\in\mathbb{R}^{l\times m}$. Let $(j+1)$-th column vector of $\mathbf{M}\in\mathbb{R}^{l\times m}$ be denoted by $\mathbf{M}_{j}\in\mathbb{R}^{l\times 1}$ so that $\mathbf{M}$ can be written as $\mathbf{M}=\left(\mathbf{M}_{0}\,|\,\cdots\,|\,\mathbf{M}_{m-1}\right)$. In particular, when $l=1$, $\mathbf{M}_{j}$ denotes the $(j+1)$-th coordinate of the row vector $\mathbf{M}$. From Lemma 1, the matrix $\mathsf{M}_{\phi}(b)$ can be written as
@@
$(\mathsf{M}_{\phi}(b)_{j})^{T}=\big{(}\mathsf{V}(b)_{j},(\mathsf{V}(b)\mathbf{P})_{j},\ldots,(\mathsf{V}(b)\mathbf{P}^{n-1})_{j}\big{)}=\big{(}\!\left<\mathsf{V}(b),((\mathbf{I}_{n})_{j}\!)^{T}\!\right>,\left<\mathsf{V}(b),(\mathbf{P}_{j}\!)^{T}\!\right>\!,\ldots,\!\left<\mathsf{V}(b),((\mathbf{P}^{n-1})_{j}\!)^{T}\!\right>\!\big{)}\enspace.$ (4)
RLWE to LWE. Let ${\bm{a}}=\mathsf{V}(a)$ for $a\in\mathcal{R}$. Define a unitary operation,
$U_{\phi,j}:\left|{\bm{a}}\,\right\rangle\left|0\,\right\rangle\mapsto\left|{\bm{a}}\,\right\rangle\left|{\bm{a}}_{j}\,\right\rangle,\quad j\in\{0,\ldots,n-1\}\enspace,$
where,
${\bm{a}}_{j}=(\mathsf{M}_{\phi}(a)_{j})^{T}=\big{(}\!\left<{\bm{a}},((\mathbf{I}_{n})_{j})^{T}\right>,\left<{\bm{a}},(\mathbf{P}_{j})^{T}\right>,\ldots,\left<{\bm{a}},((\mathbf{P}^{n-1})_{j})^{T}\right>\!\big{)}\enspace.$